# RoboTwin 2.0: A Scalable Data Generator and Benchmark with Strong Domain Randomization for Robust Bimanual Robotic Manipulation

### v1

@@ Abstract
Simulation-based data synthesis has emerged as a powerful paradigm for enhancing real-world robotic manipulation. However, existing synthetic datasets remain insufficient for robust bimanual manipulation due to two challenges: (1) the lack of an efficient, scalable data generation method for novel tasks, and (2) oversimplified simulation environments that fail to capture real-world complexity. We present RoboTwin 2.0, a scalable simulation framework that enables automated, large-scale generation of diverse and realistic data, along with unified evaluation protocols for dual-arm manipulation. We first construct RoboTwin-OD, a large-scale object library comprising 731 instances across 147 categories, each annotated with semantic and manipulation-relevant labels. Building on this foundation, we develop an expert data synthesis pipeline that combines multimodal large language models (MLLMs) with simulation-in-the-loop refinement to generate task-level execution code automatically. To improve sim-to-real transfer, RoboTwin 2.0 incorporates structured domain randomization along five axes: clutter, lighting, background, tabletop height and language instructions, thereby enhancing data diversity and policy robustness. We instantiate this framework across 50 dual-arm tasks spanning five robot embodiments, and pre-collect over 100,000 domain-randomized expert trajectories. Empirical evaluation shows a 10.9% gain in code generation success rate and improved generalization to novel real-world conditions. A vision-language-action (VLA) model fine-tuned on our dataset achieves a 367% relative improvement (42.0% vs. 9.0%) on unseen scene real-world tasks, while zero-shot models trained exclusively on our synthetic data attain a 228% relative gain, demonstrating strong generalization without real-world supervision. We release the data generator, benchmark, pre-collected dataset, and code to support scalable research in robust bimanual manipulation.
Figure 1: Overview of RoboTwin 2.0. RoboTwin 2.0 is a scalable framework for data generation and benchmarking in bimanual robotic manipulation. It integrates an expert data generation pipeline and a 50-task benchmark built on the RoboTwin Object Dataset (731 objects, 147 categories). A multimodal language model agent enables automatic task program synthesis, while flexible dual-arm configurations facilitate scalable and diverse data collection. Policies trained on RoboTwin 2.0 data demonstrate improved robustness and generalization to unseen environments.
@@ 2 Method
Figure 2: RoboTwin 2.0 Pipeline. Leveraging RoboTwin-OD and a predefined skill API, our framework employs an MLLM-based expert code generation module with simulation-in-the-loop feedback to automatically synthesize expert-level task programs. These programs are used to generate diverse, domain-randomized trajectories that support downstream policy training and evaluation.
@@ 2.1 Expert Code Generation via MLLMs and Simulation-in-the-Loop Feedback
Recent advances in language models have demonstrated their capability to generate intermediate task representations—such as textual plans hu2025text2world ; hu2024hiagent , API calls, or executable code mu2024robocodex ; roboscript —for complex robotic tasks. Multimodal large language models (MLLMs) extend this capability by incorporating visual and proprioceptive signals, enabling more grounded reasoning over real-world sensory inputs. However, most prior systems either rely on strong manual priors or lack robust closed-loop feedback during program synthesis, limiting their reliability in diverse or dynamic environments.
Building on this foundation, we propose an automated expert data generation pipeline that integrates programmatic code synthesis with multimodal execution feedback to produce high-quality manipulation programs, as illustrated in Figure 3. The system operates through a closed-loop architecture comprising two AI agents: a code-generation agent and a vision-language model (VLM) observer. By executing and monitoring code within a simulated environment, the observer systematically detects execution failures and suggests corrections, enabling the code-generation agent to iteratively refine the task program. This feedback loop facilitates the generation of robust, self-improving expert data with minimal human supervision.
@@
Multimodal Observation and Error Localization. In parallel with execution, a vision-language model (VLM) agent observes the robot’s behavior across all ten trials. The VLM performs frame-by-frame inspection to evaluate the success of each program step and localize the point of failure when errors occur. Beyond temporal localization, the VLM also diagnoses failure modes by inferring whether the underlying cause stems from flawed logic, incorrect API usage, or other systemic issues. This diagnostic capability enables the system to address root causes rather than merely responding to superficial execution errors.
Code Repair and Iterative Refinement. The code-generation agent receives two complementary feedback signals: (i) a quantitative execution log and (ii) a qualitative, localized diagnostic from the VLM. It integrates these inputs to revise the program by modifying or replacing instructions identified as failure-prone. The updated program is then re-evaluated in the next iteration. This refinement loop continues until predefined termination criteria are satisfied, yielding expert-level task code with minimal human supervision.
Termination Criteria. The refinement process terminates under one of two conditions: (i) the generated program achieves a success rate exceeding 50% across ten simulated executions in a single iteration, or (ii) the system fails to reach this threshold after five consecutive iterations. These criteria prevent indefinite refinement and ensure that only programs meeting a minimum standard of task competence are retained.
@@ 2.2 Domain Randomization for Robust Robotic Manipulation
To improve policy robustness to real-world environmental variability, we apply domain randomization across five key dimensions: (1) cluttered placement of task-irrelevant objects, (2) background textures, (3) lighting conditions, (4) tabletop heights, and (5) diverse language instructions. This systematic diversification enriches the training data distribution and significantly improves generalization to unseen scenarios. Fig. 4 (a) visualizes the effects of our domain randomization strategy.
@@
Scene Clutter. To improve policy robustness to environmental variation, we introduce cluttered tabletop scenes by randomly populating the workspace with task-irrelevant distractor objects. Leveraging our internally constructed object asset library, RoboTwin-OD, illustrated in Section 3.1, which contains 731 fully annotated objects across 147 categories, we synthesize diverse and semantically rich cluttered scenes during data generation. Each object is annotated with placement points, enabling a generic placement API that inserts arbitrary objects into the scene with semantically valid poses. To ensure physical plausibility, we precompute collision volumes and apply collision-aware placement, avoiding unrealistic overlaps introduced by randomization. Additionally, we annotate intra-class similarity groups within RoboTwin-OD to encode visual and functional similarity. During clutter sampling, we restrict the placement of distractors that are visually or semantically similar to task-relevant objects, thereby reducing ambiguity and mitigating potential policy confusion.
Diverse Background Textures. We randomize tabletop surfaces and surrounding backgrounds using a large library of textures generated via procedural methods and generative models. This exposes policies to a broad visual distribution and mitigates overfitting to clean, synthetic environments. To construct this texture library, we first leverage large language models (LLMs) combined with web crawling to collect 1,000 diverse text prompts describing real-world surface appearances, varying in style, color, and granularity. Using these prompts, we employ Stable Diffusion v2 to synthesize 20 texture samples per description, resulting in an initial pool of 20,000 candidate textures. A human-in-the-loop verification step filters out low-quality or irrelevant samples, yielding a curated set of 12,000 high-quality textures. This library is used to randomize background environments and tabletop surfaces during simulation, significantly enriching the visual diversity of the generated data. Representative samples are shown in Fig. 4 (b).
@@
Lighting Variation. Real-world environments exhibit significant lighting variability, including differences in color temperature, light source type (e.g., point lights, area lights), number of sources, and spatial configuration. Such variations alter object appearance, shading, and reflections in 2D visual inputs, posing challenges for vision-based manipulation policies. To improve robustness under diverse lighting conditions, we apply lighting randomization in our simulation pipeline. Specifically, we randomize the color, type, intensity, and position of light sources within physically plausible bounds to simulate realistic illumination diversity. As shown in Fig. 4 (second row), variations in color temperature can dramatically shift the perceived appearance of objects—for instance, changing the color of a shoe under warm versus cool lighting. By training under diverse lighting configurations, policies become more robust to real-world illumination shifts.
Tabletop Heights. In real-world settings, the height of manipulation platforms—typically tables—can vary due to differences in workspace layout and hardware configurations. These variations affect robot perception, kinematics, and interaction strategies, making policy generalization more challenging. To improve robustness to such physical discrepancies, we randomize tabletop heights during simulation. Specifically, during data generation, the table height is uniformly sampled from a physically plausible range, introducing variability in viewpoint and spatial relationships between the robot and manipulated objects.
@@
Trajectory-Level Diverse Language Instructions. To improve policy robustness to variation in natural language, we automatically generate diverse task instructions and object descriptions using a multimodal large language model. These include (i) task instruction templates and (ii) object descriptions that reflect geometry, appearance, and part-level attributes. The prompts used for generation are listed in Appendix E, with examples shown in Appendix F. For each trajectory, we construct language instructions by sampling from task instruction templates and object descriptions. For example, in the Move Can Pot task, a template such as "Use {a} to place {A} to the left of {B}" is instantiated with sampled object descriptions (e.g., {A}: "sauce can" or "white plastic lid sauce can"; {B}: "gray kitchenpot" or "kitchenpot for boiling and cooking") and robot parameters (e.g., left arm). This results in instructions such as "Use left arm to place white plastic lid sauce can to the left of kitchenpot for boiling and cooking" or "Use left arm to place sauce can to the left of gray kitchenpot." This compositional augmentation introduces linguistic variety at the trajectory level and improves generalization to unseen instructions and scene configurations.
@@ 2.3 Embodiment-Aware Grasp Adaptation
Figure 5: Five RoboTwin 2.0 Embodiment (Aloha-AgileX, ARX-X5, Piper, Franka and UR5).
@@
To address these embodiment-specific variations, we annotate each object with a diverse set of manipulation candidates that span multiple grasp axes and approach directions. This ensures that the dataset captures both manipulation diversity and robot-specific preferences. Specifically, for each object, we generate candidate grasps by incorporating preferred operation directions, randomized pose perturbations, and parallelized motion planning attempts. Additionally, we introduce angular perturbations toward directions with higher arm reachability, further expanding the space of feasible manipulation poses.
@@ 3.1 RoboTwin-OD: RoboTwin Object Dataset
For robust robotic manipulation, it is essential for policies to generalize across diverse objects, which requires a highly varied training dataset. Specifically, the dataset should include a broad range of object categories to promote general manipulation skills, as well as diverse instances within each category to capture intra-class manipulation patterns. In addition, high-quality language annotations are crucial for grounding object understanding. To this end, we developed an automated object description generator and applied it to our entire dataset, followed by human verification. For each object, we generate 15 language annotations, spanning both seen and unseen descriptions. This variation captures differences in object shape, texture, category, functionality, part structure, and description granularity.
To support learning object-centric interaction strategies, we further annotate each object with key point-axis information. These include placement points, functional points, grasping points, and grasp axis directions, explicitly encoding object affordances. Combined with our robotic manipulation API library, these annotations enable generalizable grasp execution in simulation. All objects information can be found in http://robotwin-platform.github.io/doc/objects/.
3.2 Support for Flexible Embodiment Combinations
@@
Building on our automated task generation framework, embodiment-adaptive behavior synthesis, and the large-scale object asset library RoboTwin-OD, we construct a suite of over 50 dual-arm collaborative manipulation tasks. In addition, we support data collection and evaluation across 5 distinct robot platforms, enabling comprehensive benchmarking of manipulation policies. Keyframes from representative tasks are shown in Fig. 9, and the complete task set is available at http://robotwin-platform.github.io/doc/tasks/.
@@
3.4 Pre-collected RoboTwin 2.0 Dataset
We collected over 100,000 dual-arm manipulation trajectories across 50 tasks in RoboTwin2.0, covering five distinct dual-arm embodiments. For each task-embodiment pair, we provide 100 clean (non-randomized) trajectories and 400 domain-randomized trajectories. The full dataset is available at: https://huggingface.co/datasets/TianxingChen/RoboTwin2.0/tree/main/RoboTwin2_dataset.
RoboTwin 2.0 supports a wide range of embodiment configurations and scene settings, making it a highly versatile framework for large-scale robotic data collection. It enables efficient, automated trajectory generation across diverse tasks and hardware setups with minimal human supervision.
@@ 4.1 Evaluation of Automated Expert Code Generation
In this section, we evaluate our closed-loop expert data generation system on a suite of 10 robot manipulation tasks, using programmatic generation followed by iterative refinement in simulation. Each task is specified with a natural language instruction, and for each system variant, the code-generation agent produces 10 candidate programs, each executed 10 times to account for stochasticity in dynamics, control, and perception. Task-level success is defined as the average success rate across all executions of all candidates, as described in Section 2.1.
@@
Method ASR Top5-ASR CRSR CR-Iter Token
R1.0 Vanilla 47.4% 47.4% 47.4% 1.00 1236.6
R1.0 + FB 60.4% 71.4% 60.4% 2.46 1190.4
R1.0 + MM FB 63.9% 74.2% 63.9% 2.42 1465.0
R2.0 Vanilla 62.1% 62.1% 62.1% 1.00 569.4
R2.0 + FB 66.7% 73.6% 66.7% 1.89 581.6
R2.0 + MM FB 71.3% 78.6% 71.3% 1.76 839.7
@@
Task Rate Task Rate Task Rate Task Rate
Click Alarmclock 0% Click Bell 10% Dump Bin Bigbin 0% Grab Roller 74%
Handover Block 27% Handover Mic 0% Hanging Mug 0% Lift Pot 40%
Stamp Seal 20% Turn Switch 0% Avg Success Rate 43.34%
To quantitatively characterize performance, we report five metrics: ASR (Average Success Rate), which measures mean task-level success over all candidates; Top5-ASR, the average success of the top 5 performing programs per task; CRSR (Code Revision Success Rate), the success rate after feedback-based refinement; CR-Iter, the average number of iterations taken before termination; and Token, which reflects the average LLM token cost incurred during program synthesis and repair.
Table 10 presents the evaluation results across six configurations of our system, computed over the subset of tasks supported by both RoboTwin 1.0 and 2.0 (see Appendix 8 for per-task success rates). The Vanilla setting corresponds to one-shot code generation without any iterative correction. In the FB setting, the system incorporates structured feedback from execution logs to revise failure-prone code. The MM FB configuration further augments this process with multimodal observation, allowing a vision-language model to localize and classify failures more precisely.
Across all configurations (evaluated on tasks jointly supported by RoboTwin 1.0 and 2.0), we observe that incorporating multimodal feedback consistently improves performance. For instance, within the RoboTwin 1.0 architecture, introducing feedback increases ASR from 47.4% (Vanilla) to 60.4%, and rises further to 63.9% when multimodal feedback is used. A similar trend is observed in RoboTwin 2.0, where the ASR improves from 62.1% (Vanilla) to 66.7% with execution feedback, and reaches 71.3% with the addition of vision-language-guided repair. Improvements are also evident in Top5-ASR, indicating that high-quality program candidates benefit even more from refinement strategies that incorporate perceptual understanding. These results suggest that multimodal observation not only detects failure but also provides critical context for generating more effective code repairs.
Another notable trend is the reduction in the average number of repair iterations. In the feedback-only setting, RoboTwin 2.0 requires fewer steps (1.89 vs. 2.46 in RoboTwin 1.0) to reach satisfactory performance. This pattern persists in the multimodal feedback setting, where RoboTwin 2.0 converges in 1.76 iterations on average, compared to 2.42 for RoboTwin 1.0. This suggests that RoboTwin 2.0 benefits from stronger priors in initial code generation and greater sample efficiency in refinement. Additionally, token costs are substantially reduced in RoboTwin 2.0 under the Vanilla setting (569.4 vs. 1236.6 in RoboTwin 1.0), indicating more concise and accurate initial programs.
Figure 10 visualizes the full distribution of success rates across all tasks. RoboTwin 1.0 in the Vanilla setting shows high variance and a low median, with many samples falling below the success threshold. Feedback-based configurations reduce this variance and improve central performance. In particular, RoboTwin 2.0 with multimodal feedback yields highly compact distributions centered above 80%, demonstrating both robustness and reliability. These distributional improvements reflect not only average-case performance gains but also stronger guarantees in worst-case scenarios—crucial for downstream deployment in real-world robotic systems.
Table 2 reports per-task success rates for the R2.0 + MM FB configuration across all RoboTwin 2.0-supported tasks. While many tasks yield high success (e.g., 100% on “Adjust Bottle” and “Pick Dual Bottles”), others remain unsolved. Tasks with 0% success—such as “Open Laptop” and “Shake Bottle”—highlight limitations of the code generation model in handling articulated objects, precise pose constraints, and dynamic interactions. These behaviors are difficult to specify through current action abstractions and remain challenging for program synthesis without richer multimodal memory or fine-grained control.
Taken together, these results highlight three key findings. First, vision-language feedback meaningfully improves the quality and precision of code revisions, enabling the system to address not just what failed but why it failed. Second, architectural improvements in RoboTwin 2.0 reduce both iteration count and language model token cost, enhancing efficiency. Third, the combination of symbolic execution logs with perceptual diagnostics yields expert data that is more reliable, more scalable, and more aligned with task-level semantics. These observations validate the closed-loop, self-improving architecture proposed in this work and underscore the value of multimodal reasoning in grounded code generation systems.

### v2

@@ Abstract
Simulation-based data synthesis has emerged as a powerful paradigm for enhancing real-world robotic manipulation. However, existing synthetic datasets remain insufficient for robust bimanual manipulation due to two challenges: (1) the lack of an efficient, scalable data generation method for novel tasks, and (2) oversimplified simulation environments that fail to capture real-world complexity. We present RoboTwin 2.0, a scalable simulation framework that enables automated, large-scale generation of diverse and realistic data, along with unified evaluation protocols for dual-arm manipulation. We first construct RoboTwin-OD, a large-scale object library comprising 731 instances across 147 categories, each annotated with semantic and manipulation-relevant labels. Building on this foundation, we develop an expert data synthesis pipeline that combines multimodal large language models (MLLMs) with simulation-in-the-loop refinement to generate task-level execution code automatically. To improve sim-to-real transfer, RoboTwin 2.0 incorporates structured domain randomization along five axes: clutter, lighting, background, tabletop height and language instructions, thereby enhancing data diversity and policy robustness. We instantiate this framework across 50 dual-arm tasks spanning five robot embodiments. Empirical evaluation shows a 10.9% gain in code generation success rate. Building on this, we evaluate downstream policy learning. With a mix of large-scale synthetic data and only 10 real demonstrations, a vision–language–action (VLA) model achieves a 367% relative improvement over the 10-demo baseline. Even without real data, zero-shot models trained solely on synthetic data obtain a 228% relative gain, highlighting the effectiveness of our dataset in strengthening sim-to-real transfer and robustness to environmental variations. We release the data generator, benchmark, pre-collected dataset, and code to support scalable research in robust bimanual manipulation.
Figure 1: Overview of RoboTwin 2.0. RoboTwin 2.0 is a scalable framework for bimanual manipulation, integrating an expert data generation pipeline with a 50-task benchmark built on the RoboTwin Object Dataset (731 objects, 147 categories). A multimodal language agent automates task program synthesis, while flexible dual-arm configurations enable large-scale, diverse data collection. Policies trained on RoboTwin 2.0 exhibit improved robustness and generalization to unseen environments.
@@ 2 Method
Figure 2: RoboTwin 2.0 Pipeline. Built on RoboTwin-OD and a skill API, the framework uses MLLM-based code generation with simulation feedback to produce expert task programs and domain-randomized trajectories for policy training and evaluation.
@@ 2.1 Expert Code Generation via MLLMs and Simulation-in-the-Loop Feedback
Recent advances in language models show strong ability to generate intermediate task representations—such as textual plans hu2024hiagent , API calls, or executable code mu2024robocodex ; roboscript ; hu2025text2world —for complex robotic tasks. Multimodal large language models (MLLMs) further extend this capability by incorporating visual and proprioceptive inputs, enabling more grounded reasoning over real-world sensory data. However, prior systems often depend on strong manual priors or lack closed-loop feedback during program synthesis, which limits their robustness in diverse or dynamic environments.
To address these limitations, we propose an automated expert data generation pipeline that integrates programmatic code synthesis with multimodal execution feedback (Fig.3). The system adopts a closed-loop architecture with two agents: a code-generation agent and a vision–language model (VLM) observer. The code agent synthesizes task programs from instructions, while the observer monitors execution in simulation, detects failures, and suggests corrections. This iterative feedback loop enables the code agent to refine programs automatically, producing robust, self-improving expert data with minimal human supervision. Unlike prior MLLM-based pipelines such as GenSim2huagensim2 and RoboGen wang2023robogen , our system supports zero-shot generation of complex dual-arm behaviors beyond primitive pick-and-place actions.
@@
Multimodal Observation and Error Localization. In parallel with execution, a vision-language model (VLM) agent observes the robot’s behavior across all ten trials. The VLM performs frame-by-frame inspection to evaluate the success of each program step and localize the point of failure when errors occur. Beyond temporal localization, the VLM also diagnoses failure modes by inferring whether the underlying cause stems from flawed logic, incorrect API usage, or other systemic issues. This diagnostic capability enables the system to address root causes rather than merely responding to superficial execution errors. The detail of VLM observation is shown in G.4.
Code Repair and Iterative Refinement. The code-generation agent receives two complementary feedback signals: (i) a quantitative execution log and (ii) a qualitative, localized diagnostic from the VLM. It integrates these inputs to revise the program by modifying or replacing instructions identified as failure-prone. The updated program is then re-evaluated in the next iteration, and the process continues until either the program achieves the setting success rate across ten simulated runs in one iteration or fails to do so after five consecutive refinements. This loop yields expert-level task code with minimal human supervision while avoiding indefinite refinement.
@@ 2.2 Domain Randomization for Robust Robotic Manipulation
To enhance policy robustness to real-world variability, we apply domain randomization along five dimensions: (1) cluttered distractor objects, (2) background textures, (3) lighting conditions, (4) tabletop heights, and (5) diverse language instructions. This systematic augmentation broadens the training distribution and markedly improves generalization to unseen scenarios. The effects of these randomizations are visualized in Fig. 4a.
@@
Scene Clutter. To enhance robustness to environmental variation, we augment tabletop scenes with task-irrelevant distractors drawn from RoboTwin-OD (731 objects across 147 categories; see Section 3.1). Each object includes placement annotations, enabling a generic API for semantically valid insertion. We ensure physical plausibility via collision-aware placement and precomputed volumes. To avoid policy confusion, distractors visually or semantically similar to task-relevant objects are excluded during sampling. This yields diverse yet unambiguous cluttered scenes for training.
Diverse Background Textures. We randomize tabletop surfaces and backgrounds using a large curated texture library. To build it, we first collected 1,000 diverse surface descriptions via LLM prompting and web crawling, then used Stable Diffusion v2 to generate 20 samples per description (20,000 total). After human-in-the-loop filtering, we obtained 11,000 high-quality textures. This library is applied in simulation to enrich visual diversity and reduce overfitting to clean synthetic environments (see Fig. 4b).
@@
Lighting Variation. Real-world environments exhibit diverse illumination conditions, with variations in color temperature, source type, number, and placement. These factors alter object appearance and reflections, challenging vision-based manipulation. To enhance robustness, we randomize light color, type, intensity, and position within physically plausible bounds. As shown in Fig. 4a (second row), changes in color temperature can drastically shift object appearance (e.g., a shoe under warm vs. cool light). Training under such randomized conditions improves policy robustness to real-world illumination shifts.
Tabletop Heights. In practice, table heights vary across workspaces, affecting robot perception, kinematics, and interaction. To improve generalization, we uniformly randomize table height within a plausible range during simulation, introducing variability in viewpoints and spatial relations between robot and objects.
@@
Trajectory-Level Diverse Language Instructions. To improve robustness to natural language variation, we use a multimodal LLM to generate diverse task templates and multiple object descriptions capturing geometry, appearance, and part-level attributes. Each task and object thus has several alternative phrasings, which can be flexibly combined. For every trajectory, we sample from these pools to compose instructions. For example, in Move Can Pot, the template “Use a to place A to the left of B” may yield diverse instructions such as “Use left arm to place sauce can to the left of gray kitchenpot” or “Use left arm to place white plastic lid sauce can to the left of kitchenpot for boiling and cooking.” This combinatorial augmentation produces a large set of linguistically varied instructions and significantly improves generalization to unseen language and scene configurations (see Appendix H, I).
@@ 2.3 Embodiment-Aware Grasp Adaptation
Figure 5: Five RoboTwin 2.0 Embodiments.
@@
To address embodiment-specific variations, we annotate each object with a rich set of candidate manipulation poses that cover multiple grasp axes and approach directions. This design captures both manipulation diversity and robot-specific preferences. To further expand the feasible space, we apply angular perturbations biased toward directions with higher arm reachability. Concretely, for each object we generate candidate grasps by combining preferred operation directions, randomized pose perturbations, and parallelized motion planning attempts.
@@ 3.1 RoboTwin-OD: RoboTwin Object Dataset
For robust manipulation, policies must generalize across diverse objects, which requires datasets with broad category coverage and varied intra-class instances. To facilitate language grounding, we developed an automated object description generator with human verification, producing 15 annotations per object that vary in shape, texture, functionality, part structure, and granularity.
To further support object-centric interaction, we annotate each object with key point–axis information, including placement points, functional points, grasp points, and grasp axes, explicitly encoding affordances. Together with our manipulation API library, these annotations enable generalizable grasp execution in simulation. All object information is available at http://robotwin-platform.github.io/doc/objects/ .
@@
Building on our automated task generation framework, embodiment-adaptive behavior synthesis, and the large-scale RoboTwin-OD asset library, we construct a suite of 50+ dual-arm collaborative manipulation tasks. We further support data collection and evaluation on five distinct robot platforms, enabling comprehensive cross-embodiment benchmarking. Keyframes from representative tasks are shown in Fig. 8, and the complete task descriptions are available at http://robotwin-platform.github.io/doc/tasks/. We also pre-collected over 100,000 dual-arm manipulation trajectories across 50 tasks in RoboTwin 2.0, which are available at https://huggingface.co/datasets/TianxingChen/RoboTwin2.0/tree/main/dataset.
@@ 4.1 Evaluation of Automated Expert Code Generation
We evaluate our closed-loop expert data generation system on a suite of 10 robotic manipulation tasks, each specified with a natural language instruction. For each configuration, the code-generation agent produces multiple candidate programs, which are executed in simulation to account for stochasticity in dynamics, control, and perception. Task-level success is defined as the average success rate across all executions, as described in Section 2.1.
@@
Method ASR Top5-ASR CR-Iter Token
R1.0 Vanilla 47.4% 57.6% 1.00 1236.6
R1.0 + FB 60.4% 71.4% 2.46 1190.4
R1.0 + MM FB 63.9% 74.2% 2.42 1465.0
R2.0 Vanilla 62.1% 68.0% 1.00 569.4
R2.0 + FB 66.7% 73.6% 1.89 581.6
R2.0 + MM FB 71.3% 78.6% 1.76 839.7
@@
We evaluate performance with four metrics: ASR (Average Success Rate), Top5-ASR (success over the top-5 candidates per task), CR-Iter (average refinement iterations before termination), and Token (average number of tokens in generated policy code). Results on RoboTwin 1.0 and 2.0 are reported in Table 9 under three configurations: Vanilla (one-shot code generation), FB (feedback-based repair via execution logs), and MM FB (multimodal feedback with vision–language diagnostics). Per-task success rates are provided in Appendix 8.
Across all settings, multimodal feedback yields consistent gains. In RoboTwin 1.0, ASR improves from $47.4\%$ (Vanilla) to $63.9\%$ (MM FB); in RoboTwin 2.0, it rises from $62.1\%$ to $71.3\%$. Improvements are also evident in Top5-ASR, suggesting that perceptual feedback disproportionately benefits the best candidate programs. RoboTwin 2.0 converges faster than 1.0 (e.g., $1.76$ vs. $2.42$ CR-Iter in MM FB), indicating stronger priors and more efficient refinement. Token cost is also substantially reduced, especially in Vanilla (569.4 vs. 1236.6), reflecting more concise initial code.
Figure 9 further shows that feedback narrows the success-rate distribution and raises the median. RoboTwin 2.0 with multimodal feedback achieves compact distributions centered above 80%, highlighting robustness and reliability.
Overall, three findings emerge: (1) vision–language feedback not only detects failures but also guides precise repairs; (2) architectural improvements in RoboTwin 2.0 accelerate convergence and reduce token usage; and (3) combining symbolic execution logs with perceptual diagnostics yields more reliable, semantically aligned expert data. Together, these results validate the effectiveness of our closed-loop, self-improving code generation architecture. Detailed setups, metric definitions, and additional analyses are provided in Appendix G.